40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 12 08:40:06 2016, Steve, Update, PEM, Air cond maintenance 
    Reply  Mon Jan 25 11:06:08 2016, Steve, Update, PEM, noisy fan belt replaced 
    Reply  Fri Jan 29 08:18:13 2016, Steve, Update, PEM, PEM PEM.png
       Reply  Wed Apr 20 14:36:45 2016, Steve, Update, PEM, dusty lab AP_accm.jpgwiped_AP.jpgtemp-partic_6months.png3_years_PEM.png
          Reply  Tue Jun 21 14:12:10 2016, Steve, Update, PEM, vacuum envelope wiped 60dPEM.png
       Reply  Fri Jan 13 14:18:28 2017, Steve, Update, PEM, air condition fixed PEM_120d.png
Message ID: 12209     Entry time: Tue Jun 21 14:12:10 2016     In reply to: 12084
Author: Steve 
Type: Update 
Category: PEM 
Subject: vacuum envelope wiped 

I wiped down the cranes with wet towel and Mario our janitor did the chamber tops with the tubes.

The optical tabels were not touched.

The outside temp peaked at 44 C yesterday

 

Attachment 1: 60dPEM.png  56 kB  | Hide | Hide all
60dPEM.png
ELOG V3.1.3-